OpenStudy (anonymous):

A circular arc of length 6 ft subtends a central angle of 65 degrees. Find the radius of the circle in ft.

OpenStudy (mathmale):

How is arc length related to the radius and to the "subtended angle" (or "central angle") in a circle?

OpenStudy (anonymous):

Length of circular arc = (angle in radians)*(radius of the circle) An obvious example is the circumference of a circle given by 2*pi*R (2*pi being 360 degrees) So call the angle A and we have Length of arc = A*R

OpenStudy (anonymous):

In the here and now the angle is 65 degrees which is 65*pi/180 radians

OpenStudy (anonymous):

\[65*\pi/180 radians=1.1344640138\]

OpenStudy (anonymous):

(I use this to convert from degrees to radians and vice versa since there are pi radians in 180 degrees)

OpenStudy (anonymous):

R = (length of arc)/A = 6/1.1344640138 = 5.28885925492 ft

OpenStudy (anonymous):

Do you get it or no?

OpenStudy (anonymous):

yes thank youuu!

OpenStudy (anonymous):

yw next time I'd like it if you'd try and put a little input in so that I can "help" you instead of just give you the answer